two-thirds of 5/6 of a number x is the same as 3/4 of 4/5 of a number y. what is the value of x/y as a fraction in lowest terms?
Answer by user_dude2008(1825)

(2/3)(5/6)x=(3/4)(4/5)y


(10/18)x=(12/20)y


(5/9)x=(3/5)y

(5/3)(5/9)x=(5/3)(3/5)y


(25/27)x=y


y = (25/27)x



---------> x/y = x/[(25/27)x] = 27/25

Answer: x/y = 27/25